Operations Report – March 1st

Crew 206 Operations Report 01 – MAR – 2019

SOL: 5

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 10,5 hours

Solar— SOC % – Turned off Gen : 100%, Turned on Gen : 69%

Diesel Reading – 75%

Propane Reading – 85%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 470 gallons

Water differential (Static tank) – 10 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 264 gallons

Static to Loft Pump used – yes

Water Meter: 0141620,0 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not used

Hours: 53,0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used

Hours: 77,3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,4 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Still on the fixing with BDM CapCom.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Noticed a disconnected tarp on
the connector between the Green Hab and the Science Dome
during this morning’s EVA. We have replaced it rapidly,
but it still needs fixation, especially considering the
wind. We will tackle this issue on the next EVA.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.